OpenStudy (anonymous):

What is a factored form of 2x^2-6x ? Please explain! I will become a fan of anyone who answers!

OpenStudy (anonymous):

each term has a common factor of \(2x\)

OpenStudy (anonymous):

\[2x^2=2x\times x\] and \[-6x=2x\times (-3)\]so you can factor it as \[2x(x-3)\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw